COLD CHEMICAL REACTIONS OF H-ATOMS AND N_{2}O IN SOLID PARAHYDROGEN AT 2 K

\underline{DAVID T. ANDERSON} AND FREDRICK M. MUTUNGA, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

We report on the results of chemical reactions of H-atoms with N_{2}O in solid parahydrogen matrices investigated by rapid scan FTIR. The reaction is initiated by the 193 nm in situ photodissociation of N_{2}O to produce N_{2} + O(^{1}D). We assume the O(^{1}D) photoproduct reacts with the parahydrogen host to create H-atoms. What we observe is growth of cis-HNNO right after photolysis is stopped along with a later delayed growth in trans-HNNO. We assign our peaks by comparison with a previous Xe matrix isolation study^{a} and recent ab initio calculations^{b} of the anharmonic vibrational frequencies and isotopic shifts. We performed some experiments using the ^{15}N_{2}^{18}O precursor molecule. Detailed study of the reaction kinetics indicate that at 1.8 K the cis-isomer is formed exclusively by the reaction, but then converts to the lower energy trans-isomer with time. These results are surprising because a 40 kJ/mol barrier is predicted for the H + N_{2}O reaction, yet the reaction readily proceeds even at 1.8 K. We are currently studying this reaction under a variety of conditions and the most current results and analysis will be presented.

aS. L. Laursen, A. E. Delia, and K. Mitchell, J. Phys. Chem. A 104, 3681-3692 (2000).

bK. A. Petersen and J. S. Francisco, J. Chem. Phys. 134, 084308 (2011).